OpenStudy (anonymous):

What is the area of a circle with a radius of 8? If necessary, round your answer to two decimal places. Everytime i solve these problems i somehow get the wrong answer so how do i solve this so i can see if im even doing it remotely right

OpenStudy (anonymous):

A=pi*r^2 A=pi*64=201.06

OpenStudy (anonymous):

oh so i was doing it right i must be putting my answers in wrong XD thank you!